Given the reactions below, answer the following questions.

Cl2(g) + F2(g) = 2ClF(g) \(\Delta G_{\text{rxn}}^\circ\) = 115.4 kJ/mol

Cl2(g) + Br2 (g) =2ClBr(g) $$  \(\Delta G_{\text{rxn}}^\circ\) = –2.0 kJ/mol


Calculate the Δ?∘rxn for

2ClF(g) +Br2(g) = 2ClBr(g) + F2(g)
Answer #1

Cl2(g) + F2(g) = 2ClF(g)    Δ?∘ = 115.4 kJ/mol

Cl2(g) + Br2 (g) =2ClBr(g) Δ?∘ = –2.0 kJ/mol

Cl2(g) + Br2 (g) =2ClBr(g) Δ?∘ = –2.0 kJ/mol

Cl2(g) + F2(g) = 2ClF(g)    Δ?∘ = 115.4 kJ/mol       subtract two equations

(-)        (-)          (-)                     (-)

--------------------------------------------------------------------------------------------------

2ClF(g) +Br2(g) -------->2ClBr(g) + F2(g)     Δ?∘rxn   = -117.4KJ/mole    >>>>>answer